OpenStudy (anonymous):

The sum of three consecutive even integers is -72. What is the least of these numbers? -26 -22 -30 24

OpenStudy (anonymous):

Are you sure those are the answers?

OpenStudy (campbell_st):

let x be 1 even integer what is the next one..?

OpenStudy (anonymous):

Nevermind I read it wrong

OpenStudy (denonakavro):

Think as the consecutive numbers as this (x+1)+(x+2)+(x+3)= -72 Solve the equation and then find x So would look like this 3x+6=-72 3x=-78 x= -78/3

OpenStudy (denonakavro):

Read my previous post and x would be your answer

OpenStudy (anonymous):

Because the consecutive numbers have to be even so you would have to do x + (x + 2) + (x + 4) because then you would be getting all consecutive even numbers.

OpenStudy (denonakavro):

Sorry I am wrong they are supposed to be even listen to Lexie

OpenStudy (campbell_st):

are you sure x is the answer... remember you have x, x+2 and x + 4

OpenStudy (anonymous):

Isn't "least" referring to greatest negative value?

OpenStudy (triciaal):

@denonakavro we are doing even consecutive so x is even and count by 2's ( x-2, x, x + 2 )

OpenStudy (anonymous):

x + (x +2) + (x + 4) = -72; afterwards you would add all the xs together and add the numbers and get 3x + 6 = -72 then afterwards you would subtract 6 from both sides and end up with 3x = -78 then you would divide 3 on both sides and end with x = -26 but that is the greatest factor like @jaeuni had said so you would have to take -26 and add 4 to it

OpenStudy (triciaal):